OpenStudy (anonymous):

The corresponding sides of two similar rectangles are 14 ft and 8ft. The area of the smaller rectangle is 57ft^2. To the nearest whole number , what is the area of the smaller rectangle?

OpenStudy (anonymous):

area are in ratio 8^2 : 14^2 or 64 / 196

OpenStudy (anonymous):

so we have 64 / 196 = 57/ a a =( 57 * 196 )/ 64

OpenStudy (anonymous):

you want the area of larger rectangle right?

OpenStudy (anonymous):

if so it is 174.56

OpenStudy (anonymous):

ok? you said u wanted area of smaller but u gave area of smaller so i assume larger rectangle

OpenStudy (anonymous):

yes i want the larger sorry typo

OpenStudy (anonymous):

thats ok

OpenStudy (anonymous):

do u follow the calculation ok?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

alright - plz click good answer box when u are happy with it so as to close out the question thnx